OpenStudy (anonymous):

For each of the following find lim h->0 [f(x+h)-f(x)]/h f(x)=x^2+1 f(x)=x^2+2x+1 f(x)= x^3

OpenStudy (mertsj):

\[\frac{(x+h)^2+1-(x^2+1)}{h}=\frac{x^2+2xh+h^2+1-x^2-1}{h}\]

OpenStudy (anonymous):

I'll work the first one for you. The others are up to you. \[f(x)=x^2+1\] \[\frac{f(x+h)-f(x)}{h}=\frac{(x+h)^2+1-x^2-1}{h}\] \[=\frac{x^2+2xh+h^2+1-x^2-1}{h}\] \[=\frac{2xh+h^2}{h}=\frac{h(2x+h)}{h}=2x+h\] \[\lim_{h \rightarrow 0}(2x+h)=\lim_{h \rightarrow 0}(2x+0)=2x\]

OpenStudy (anonymous):

which equals 2x right for the first one

OpenStudy (mertsj):

\[\frac{2xh+h^2}{h}=2x+h\] As h approaches 0, the limit is 2x

OpenStudy (anonymous):

the second one should be h+2x+2 which by the limit would end up being 2x+2

OpenStudy (anonymous):

yes, you will learn a quicker method shortly

OpenStudy (anonymous):

i am a bit stuck with f(x)=√(x+1)